OpenStudy (help_needed_cudi):

Can anyone help me out with this problem? Just need steps! and how to do it....√8 ∙ 0.25√6

OpenStudy (anonymous):

76

OpenStudy (anonymous):

first multiple 8 with 0.25 then minus6 then add 0.43 the divide it all by 96

OpenStudy (help_needed_cudi):

please i will reward you!!!!

OpenStudy (anonymous):

are u still looking for answers here?

OpenStudy (anonymous):

@Help_needed_cudi ?

OpenStudy (mathmale):

\[.√8 ∙ 0.25√6\] Is what I think you have presented. Please confirm this, or fix it.

OpenStudy (mathmale):

\[.√8 ∙ 0.25√6\] could be re-written as \[0.25*\sqrt{8}*\sqrt{6}\] ... and \[\sqrt{8}*\sqrt{6}\] can be simplified by writing it as a single radical. Try this, please.

OpenStudy (help_needed_cudi):

oh thamks guys!! @mathmale @jeedeee @kx2bay i had figured it but but thatnks i wish i could give u each a medal